You are required to mix up 5000 US gals 10% w/w of sulphuric acid solution
(a) How much concentrated sulphuric acid do you need (US gals)
How much Water (US gals)
(c) You want to decant the concentrated acid from drums into a tank, when first setting up on location so that you can pump it out using flowmeters ‘on the fly’.
The yard has a 4000 gal fibreglass-lined acid tank or a 10 bbl carbon steel tank available …. or you may have to rent a 2000 gal stainless steel vessel from the competition at a premium rate.
What tank should you use.
Data: Conc sulphuric acid available is 98% strength, with an SG of 1.84.
10% sulphuric acid solution has a SG of 1.05.

Answer by Alan3354(69443)

You are required to mix up 5000 US gals 10% w/w of sulphuric acid solution
(a) How much concentrated sulphuric acid do you need (US gals)
How much Water (US gals)
------------
Mixing 98% acid with water to give 5000 gals of 10%:
s + w = 5000
98s + 0w = 5000*10
s = 50000/98 gallons
s =~ 510.2 gallons of acid
w = 5000 - s = 4489.8 gallons of water
===================================
(c) You want to decant the concentrated acid from drums into a tank, when first setting up on location so that you can pump it out using flowmeters ‘on the fly’.
The yard has a 4000 gal fibreglass-lined acid tank or a 10 bbl carbon steel tank available …. or you may have to rent a 2000 gal stainless steel vessel from the competition at a premium rate.
What tank should you use.
Based on what? Use the one that's most convenient.
---------------
You didn't give the volume of the 10 bbl tank.
The oil industry uses 42 gallons/bbl
Wine is 33 gal/bbl
Common barrels are 55 gallon drums.
=========================
Data: Conc sulphuric acid available is 98% strength, with an SG of 1.84.
10% sulphuric acid solution has a SG of 1.05.
------------
The SG's are not relevant.